OpenStudy (anonymous):

solving linear equations need help solving step by step 1/3(m-27)=4

OpenStudy (anonymous):

@madhu.mukherjee.946

OpenStudy (anonymous):

@Zale101

Vocaloid (vocaloid):

first step : multiply both sides by 3

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

3m and 81 but woundent it be like this 27/1 1/3 then multiply?

OpenStudy (anonymous):

then it would be simplified then 9 right

Vocaloid (vocaloid):

uh... not quite... (1/3)*3 = ?

OpenStudy (anonymous):

1/3 3/1=3/3

Vocaloid (vocaloid):

right, and 3/3 = 1, right? so (1/3)(m-27) times 3 = 1*(m-27) = m - 27, with me so far?

OpenStudy (anonymous):

yea

Vocaloid (vocaloid):

now multiply the right side by 3

Vocaloid (vocaloid):

4*3 = ?

OpenStudy (anonymous):

12

Vocaloid (vocaloid):

so now we have m - 27 = 12 add 27 to both sides

OpenStudy (anonymous):

39 M=39 holly crap thats right but i have been taught another way

OpenStudy (anonymous):

were did you get 3??

OpenStudy (anonymous):

for the first step?

Vocaloid (vocaloid):

well, we started off with 1/3(m-27), right? I want to "get rid of" (1/3) so I multiply by 3 to change 1/3(m-27) to (m-37)

OpenStudy (anonymous):

ohhh ok thanks.